OpenStudy (anonymous):

Can some please explain how to do this pre-calc question? For the function defined by: f(x)={ x^2, x≤1, 2x+1, x>1} evaluate f(0)

OpenStudy (anonymous):

as u see 0 is less than 1 Which Condition Will Suit

OpenStudy (anonymous):

x^2, x≤1 or 2x+1, x>1

OpenStudy (anonymous):

probabbly x^2, x≤1 f(0) = 0^2 = 0

OpenStudy (anonymous):

ok thanks! I got that answer too, just wanted to make sure I was doing it right!:)

OpenStudy (anonymous):

Welcome